# references.bib

### v1

@@ 5 Experiments
GAMA 7B 41.44 45.40 32.33 30.83 18.91 19.21 30.90 31.81
Qwen Audio [chuQwenAudioAdvancingUniversal2023] 55.25 56.73 44.00 40.90 30.03 27.95 43.10 41.86
Qwen2 Audio [chuQwen2AudioTechnicalReport2024a] 54.95 45.90 50.98 53.26 42.04 45.90 49.20 52.50
Mellow [deshmukhMellowSmallAudio2025] 61.26 64.90 54.19 52.67 29.73 38.77 48.40 52.11
Gemini Pro v1.5 56.75 54.46 49.40 48.56 58.55 55.90 54.90 52.97
Audio Flamingo 2 [ghoshAudioFlamingo22025] 61.56 65.10 73.95 72.90 30.93 40.26 55.48 59.42
Kimi-Audio [kimiteamKimiAudioTechnicalReport2025] 61.68 - 73.27 - 60.66 - 65.00 -
@@
Audio-Reasoner [xieAudioReasonerImprovingReasoning2025] 60.06 - 64.30 - 60.70 - 61.71 -
Audio-CoT [maAudioCoTExploringChainofThought2025] 61.86 - 56.29 - 55.26 - 57.80 -
R1-AQA [liReinforcementLearningOutperforms2025] 68.77 69.76 64.37 61.40 63.66 62.70 65.60 64.36
Qwen2.5-Omni-7B [xuQwen25OmniTechnicalReport2025] 67.87 - 69.16 - 59.76 - 65.60 -
@@ 5.1 Experimental Setup
  • MMAU (Massive Multi-Task Audio Understanding) [sakshiMMAUMassiveMultiTask2024] (Table 3): Evaluates expert-level knowledge and complex reasoning across speech, environmental sounds and music across 27 diverse tasks. Scores are reported on both Test-Mini (1k clips) and Test (10k clips) sets; some authors only report Test-Mini. MMAU is a multiple choice question benchmark, where the model has to select the answer out of four possible answers. MMAU uses the accuracy metric, as scores are being calculated by the match between the predicted answer and the correct answer. The answer part is extracted out of the model’s full response before evaluating it on MMAU.
  • AudioBench [wangAudioBenchUniversalBenchmark2025] (Table 4): A universal benchmark covering a diverse set of audio tasks, focusing on instruction following capabilities. For our evaluation, we utilize a relevant subset focusing on Audio Question Answering (AQA), including AudioCaps QA [kimAudioCapsGeneratingCaptions2019], Clotho AQA [lippingClothoAQACrowdsourcedDataset2022] and WavCaps AQA [meiWavCapsChatGPTassistedWeaklylabelled2024]. Automated Audio Captioning (AAC), including AudioCaps [kimAudioCapsGeneratingCaptions2019] and WavCaps [meiWavCapsChatGPTassistedWeaklylabelled2024], and Music Question Answering Music (QA), including MuchoMusic [weckMuChoMusicEvaluatingMusic2024]. Qwen2.5-Omni is included in the benchmarks for comparison. The AQA and Music QA benchmarks are open question answering benchmarks, in contrast to MMAU. For AQA, Music QA and AAC we evaluate the full output of the model against a model-as-judge. A Llama 3 model [llama3teamLlama3Herd2024] is tasked with evaluating the response against the correct answer.
@@
Qwen-Audio [chuQwenAudioAdvancingUniversal2023] 50.22 61.93 42.70 59.06 47.04 32.94
MERaLiON [heMERaLiONAudioLLMBridgingAudio2025] 49.78 63.15 46.32 57.79 38.00 33.98
Qwen2-Audio [chuQwen2AudioTechnicalReport2024a] 45.75 50.92 44.47 71.61 40.78 33.78
Phi4 [abdinPhi4TechnicalReport2024] 38.47 47.87 35.13 54.42 26.39 21.88
SeaLLMs [nguyenSeaLLMsLargeLanguage2024] 53.74 53.04 42.11 63.18 53.21 -
WavLLM [huWavLLMRobustAdaptive2024] 29.84 43.01 26.25 44.31 5.50 6.90
SALMONN [tangSALMONNGenericHearing2024] 50.29 57.75 47.30 50.88 37.45 23.77
Qwen2.5-Omni [xuQwen25OmniTechnicalReport2025] 59.62 59.91 53.29 70.09 51.97 41.48
@@ 5.2 Main Results
Length constraint reward impacts performance. Experiments investigating thinking budget constraints, detailed in Appendix D, demonstrate the effectiveness of the GRPO length constraint reward function (Equation 1, using default parameters $\alpha=0.1$ and $\delta=0.5$). The best results on the MMAU Test-Mini benchmark were observed with thinking budgets (25 words, and 100-150 words, see Table 6) that allowed the model to operate within a range where the reward signal was active. Excessively long target reasoning phases, if they fall outside this effective reward range, do not necessarily translate to better performance under the current setup.
@@ Appendix A Acknowledgement
Appendix B Ethics Statement
B.1 Responsible Data Science
@@ B.2 Broader Impact
Appendix C Detailed Data Generation Pipeline
C.1 Dataset Creation
@@
A second classification pass uses the Mixtral-8x7B-Instruct model [jiangMixtralExperts2024] in a zero-shot setting to verify the sound descriptions. The model is prompted to identify whether captions represent genuine audio events based on specific criteria including sound-related verbs, impact descriptions, and audio property descriptions (see prompt in Appendix F.1). Next to a Mixtral-based model, a small sample of 1000 examples of closed caption subtitles from movies and tv shows are utilized to train a simple BERT sentence classification model, which is used to filter out the captions as well. Only captions that are classified as sound descriptions by both models are kept. The remaining dataset contained 9,464,882 captions.
@@ C.2 Dataset Processing
    Qwen2Audio-7B-Instruct [chuQwen2AudioTechnicalReport2024a], a large language model fine-tuned for audio understanding, generates detailed natural language descriptions of the audio content. The model processes the 16kHz audio waveform along with a prompt requesting detailed audio description. The model outputs are decoded using beam search with a maximum length of 1024 tokens.
@@
    BEATs (Bidirectional Encoder representation from Audio Transformers) [chenBEATsAudioPretraining2023], pretrained on AudioSet-2M, performs multi-label audio tagging. The model analyzes the padded audio waveforms and outputs probability scores for audio events. The top five predicted tags and their confidence scores are retained for each segment.
@@
    AST (Audio Spectrogram Transformer) [gongASTAudioSpectrogram2021], fine-tuned on AudioSet and ESC-50, provides audio event classification. The model processes audio spectrograms and classifies them into predefined categories from the AudioSet ontology. This provides an additional layer of audio event recognition.
@@
    CoNeTTE [labbeCoNeTTEEfficientAudio2024] model analyzes the 32kHz audio to generate contextual audio event descriptions. The model leverages contrastive learning to map audio segments to semantic descriptions of audio events.
@@
    LP-MusicCaps Model [dohLPMusicCapsLLMBasedPseudo2023], a specialized BART-based music captioning model, trained on the LP-MusicCaps dataset, is employed specifically for segments containing music. The model generates natural language descriptions of musical characteristics when music is detected in the audio.
@@
    BLIP [liBLIPBootstrappingLanguageImage2022], a vision-language model, generates natural language captions for each frame. The model uses beam search with four beams and a maximum length of 20 tokens to generate descriptive captions.
@@
    CLIP [radfordLearningTransferableVisual2021] performs zero-shot classification against a predefined set of object categories. Each frame is evaluated against text prompts in the format “a photo of a [category]”, with the model’s logits being converted to probabilities through softmax normalization.
@@
    RT-DETR [zhaoDETRsBeatYOLOs2024a], a real-time object detection model, identifies objects in each frame with a confidence threshold of 0.3. The model’s outputs are post-processed to account for varying image sizes while maintaining accurate bounding box coordinates.
@@
    A Places365-trained ResNet-50 [zhouLearningDeepFeatures2014] classifies the scene/environment in each frame. The frames are resized to 256x256, center-cropped to 224x224, and normalized using ImageNet statistics before being processed by the model.
@@
C.3 Dataset Generation
The dataset generation process combines outputs from the three modality-specific pipelines into a structured format. Initial evaluations of several large language models (Qwen 2.5 [yangQwen25TechnicalReport2025], QwQ [qwenteamQwQReflectDeeply2024], Llama 3.3 [llama3teamLlama3Herd2024], Mistral-Small 3 [mistralteamMistralSmall3] and Deepseek-based models R1 Llama and R1 Qwen [guoDeepSeekR1IncentivizingReasoning2025]) show that Qwen 2.5 performs significantly better, particularly in maintaining consistent JSON output formats. Based on these results, the final dataset generation pipeline uses Qwen2.5-72B-Instruct, implementing structured generation with a schema through the xgrammar library [dongXGrammarFlexibleEfficient2025a] with the help of vLLM [kwonEfficientMemoryManagement2023]. This model with a relatively high number of parameters was chosen due to the tendency that larger models often function as teacher models for smaller student models with knowledge distillation and compression [huTeacherStudentArchitectureKnowledge2023]. The generation process follows a two-phase prompt and three-phase reasoning: a thinking phase for reasoning, a semantic descriptor phase for the key semantic descriptors, and an answer phase for the final caption, both enforced through structured JSON output (see Appendix F.2).
@@ C.3 Dataset Generation
Appendix D Thinking Budget
@@
$r(y,y_{\text{gold}},n_{\text{gold}})$ =$=\begin{cases}\text{clip}(1-\alpha(n_{\text{gold}}-n_{y})+\delta,0,1),&\text{if }n_{y}\leq n_{\text{gold}}\\ \\ \text{clip}(\alpha(n_{\text{gold}}-n_{y})+\delta,0,1),&\text{otherwise}\end{cases}$ (1)
@@
$01-\alpha(n_{gold}-n_{y})+\delta=1-0.1\cdot 15+0.5=1-1.5+0.5=$ 0 (4)
For the case where the generated thinking length exceeds the target ($n_{y}>n_{gold}$), which corresponds to the original LPCO Max Length Constraint Mode as defined in [aggarwalL1ControllingHow2025], the reward function becomes:
@@
This clipping behavior has implications for training when $\alpha$ is 0.1: if the model’s average generated thinking length is significantly shorter (e.g., more than 15 words below $n_{\text{gold}}$) than high target lengths like 125 or 150 words, it may consistently receive a zero reward component from this length constraint. This could hinder convergence towards such longer budgets, as the model does not receive a continuous gradient “nudge” from the length reward to increase its thinking length into the desired range. While adjusting $\alpha$ to a much lower value could widen the range before clipping, an overly small $\alpha$ might diminish the length reward’s impact, potentially allowing it to be overshadowed by other rewards in the GRPO objective. In experiments, decreasing $\alpha$ was found to be non-effective. It is crucial that the reward function has a slope (before clipping) to ensure effective gradient propagation in GRPO. The referenced Max Length Constraint Mode [aggarwalL1ControllingHow2025] is designed to apply a soft constraint, gradually penalizing outputs exceeding the target length rather than imposing a hard cutoff, and it incentivizes token efficiency without sacrificing correctness. The $\delta=0.5$ term helps ensure that correct answers with minor budget violations are still preferred over incorrect ones.
@@
Appendix E Example Model Outputs
@@ 1. Sound-generating animated beings with descriptive adjectives:
Appendix F Prompts
F.1 Filtering Prompt
Prompt for Caption Filtering You are a friendly chatbot whose task it is to filter out bad data. You will get a closed caption corresponding to a video clip. Your task is to state whether the caption is a correct subtitle for deaf or hard-of-hearing people. Correct captions in this task are those that correspond to words that could represent an actual sound being made. This could either include a verb that states an impact or sound types or properties like "sound", "noise" or "music". Incorrect closed captions include sentences that someone is saying in the video clip, or sentences that are not related to the video clip at all. All captions are in English. All captions are within curly brackets or square brackets []. Examples of correct captions include: - "(laughs)" or "(laughter)" - "[XBOX SOUND]" - "[chicken bocking imitation]" - "(cereal grains smacking onto wood)" - "(collision)" Examples of incorrect captions include: - "[ transport ]" - "(Wishes are left to wither by time.)" - "(look, I like my nightmareless sleep; I’ll play some scary games when I feel too peaceful)" - "[A calm navy color] [TinyTAN character detail]" - "[Haotian Sword Tower]" Is the following caption correct? Please only answer "yes" or "no"
F.2 Caption Generation Prompt
@@
F.3 Prepended Prompt on Model Generation

### v2

@@ 5 Experiments
GAMA 7B [22] 41.44 45.40 32.33 30.83 18.91 19.21 30.90 31.81
Qwen Audio [7] 55.25 56.73 44.00 40.90 30.03 27.95 43.10 41.86
Qwen2 Audio [8] 54.95 45.90 50.98 53.26 42.04 45.90 49.20 52.50
Mellow [12] 61.26 64.90 54.19 52.67 29.73 38.77 48.40 52.11
Gemini Pro v1.5 [19] 56.75 54.46 49.40 48.56 58.55 55.90 54.90 52.97
Audio Flamingo 2 [21] 61.56 65.10 73.95 72.90 30.93 40.26 55.48 59.42
Gemini 2.0 Flash [50] 56.46 61.73 58.68 56.53 51.65 61.53 55.60 59.93
Kimi-Audio [34] 61.68 - 73.27 - 60.66 - 65.00 -
@@
Audio-Reasoner [65] 60.06 - 64.30 - 60.70 - 61.71 -
Audio-CoT [44] 61.86 - 56.29 - 55.26 - 57.80 -
R1-AQA [38] 68.77 69.76 64.37 61.40 63.66 62.70 65.60 64.36
Gemini 2.5 Flash [10] 73.87 - 65.57 - 62.16 - 67.18 -
Qwen2.5-Omni-7B [66] 67.87 - 69.16 - 59.76 - 65.60 -
@@ 5.1 Experimental Setup
  • MMAU (Massive Multi-Task Audio Understanding) [54] (Table 3): Evaluates expert-level knowledge and complex reasoning across speech, environmental sounds and music across 27 diverse tasks. Scores are reported on both Test-Mini (1k clips) and Test (10k clips) sets; some authors only report Test-Mini. MMAU is a multiple choice question benchmark, where the model has to select the answer out of four possible answers. MMAU uses the accuracy metric, as scores are being calculated by the match between the predicted answer and the correct answer. The answer part is extracted out of the model’s full response before evaluating it on MMAU.
  • AudioBench [59] (Table 4): A universal benchmark covering a diverse set of audio tasks, focusing on instruction following capabilities. For our evaluation, we utilize a relevant subset focusing on Audio Question Answering (AQA), including AudioCaps QA [33], Clotho AQA [41] and WavCaps AQA [45]. Automated Audio Captioning (AAC), including AudioCaps [33] and WavCaps [45], and Music Question Answering Music (QA), including MuchoMusic [61]. Qwen2.5-Omni is included in the benchmarks for comparison. The AQA and Music QA benchmarks are open question answering benchmarks, in contrast to MMAU. For AQA, Music QA and AAC we evaluate the full output of the model against a model-as-judge. A Llama 3 model [42] is tasked with evaluating the response against the correct answer.
@@
Qwen-Audio [7] 50.22 61.93 42.70 59.06 47.04 32.94
MERaLiON [27] 49.78 63.15 46.32 57.79 38.00 33.98
Qwen2-Audio [8] 45.75 50.92 44.47 71.61 40.78 33.78
Phi4 [1] 38.47 47.87 35.13 54.42 26.39 21.88
SeaLLMs [48] 53.74 53.04 42.11 63.18 53.21 -
WavLLM [31] 29.84 43.01 26.25 44.31 5.50 6.90
SALMONN [58] 50.29 57.75 47.30 50.88 37.45 23.77
Qwen2.5-Omni [66] 59.62 59.91 53.29 70.09 51.97 41.48
@@ 5.2 Main Results
Length constraint reward impacts performance. Experiments investigating thinking budget constraints, detailed in Appendix C, demonstrate the effectiveness of the GRPO length constraint reward function (Equation 1, using default parameters $\alpha=0.1$ and $\delta=0.5$). The best results on the MMAU Test-Mini benchmark were observed with thinking budgets (25 words, and 100-150 words, see Table 6) that allowed the model to operate within a range where the reward signal was active. Excessively long target reasoning phases, if they fall outside this effective reward range, do not necessarily translate to better performance under the current setup.
@@ 7 Acknowledgement
Appendix A Ethics Statement
A.1 Responsible Data Science
@@ A.2 Broader Impact
Appendix B Detailed Data Generation Pipeline
B.1 Dataset Creation
@@
A second classification pass uses the Mixtral-8x7B-Instruct model [32] in a zero-shot setting to verify the sound descriptions. The model is prompted to identify whether captions represent genuine audio events based on specific criteria including sound-related verbs, impact descriptions, and audio property descriptions (see prompt in Appendix E.1). Next to a Mixtral-based model, a small sample of 1000 examples of closed caption subtitles from movies and tv shows are utilized to train a simple BERT sentence classification model, which is used to filter out the captions as well. Only captions that are classified as sound descriptions by both models are kept. The remaining dataset contained 9,464,882 captions.
@@ B.2 Dataset Processing
    Qwen2Audio-7B-Instruct [8], a large language model fine-tuned for audio understanding, generates detailed natural language descriptions of the audio content. The model processes the 16kHz audio waveform along with a prompt requesting detailed audio description. The model outputs are decoded using beam search with a maximum length of 1024 tokens.
@@
    BEATs (Bidirectional Encoder representation from Audio Transformers) [6], pretrained on AudioSet-2M, performs multi-label audio tagging. The model analyzes the padded audio waveforms and outputs probability scores for audio events. The top five predicted tags and their confidence scores are retained for each segment.
@@
    AST (Audio Spectrogram Transformer) [24], fine-tuned on AudioSet and ESC-50, provides audio event classification. The model processes audio spectrograms and classifies them into predefined categories from the AudioSet ontology. This provides an additional layer of audio event recognition.
@@
    CoNeTTE [37] model analyzes the 32kHz audio to generate contextual audio event descriptions. The model leverages contrastive learning to map audio segments to semantic descriptions of audio events.
@@
    LP-MusicCaps Model [14], a specialized BART-based music captioning model, trained on the LP-MusicCaps dataset, is employed specifically for segments containing music. The model generates natural language descriptions of musical characteristics when music is detected in the audio.
@@
    BLIP [40], a vision-language model, generates natural language captions for each frame. The model uses beam search with four beams and a maximum length of 20 tokens to generate descriptive captions.
@@
    CLIP [52] performs zero-shot classification against a predefined set of object categories. Each frame is evaluated against text prompts in the format “a photo of a [category]”, with the model’s logits being converted to probabilities through softmax normalization.
@@
    RT-DETR [70], a real-time object detection model, identifies objects in each frame with a confidence threshold of 0.3. The model’s outputs are post-processed to account for varying image sizes while maintaining accurate bounding box coordinates.
@@
    A Places365-trained ResNet-50 [71] classifies the scene/environment in each frame. The frames are resized to 256x256, center-cropped to 224x224, and normalized using ImageNet statistics before being processed by the model.
@@
B.3 Dataset Generation
The dataset generation process combines outputs from the three modality-specific pipelines into a structured format. Initial evaluations of several large language models (Qwen 2.5 [67], QwQ [51], Llama 3.3 [42], Mistral-Small 3 [46] and Deepseek-based models R1 Llama and R1 Qwen [25]) show that Qwen 2.5 performs significantly better, particularly in maintaining consistent JSON output formats. Based on these results, the final dataset generation pipeline uses Qwen2.5-72B-Instruct, implementing structured generation with a schema through the xgrammar library [15] with the help of vLLM [36]. This model with a relatively high number of parameters was chosen due to the tendency that larger models often function as teacher models for smaller student models with knowledge distillation and compression [29]. The generation process follows a two-phase prompt and three-phase reasoning: a thinking phase for reasoning, a semantic descriptor phase for the key semantic descriptors, and an answer phase for the final caption, both enforced through structured JSON output (see Appendix E.2).
@@ B.3 Dataset Generation
Appendix C Thinking Budget
@@
[…]
@@
$1-\alpha(n_{gold}-n_{y})+\delta=1-0.1\cdot 15+0.5=1-1.5+0.5=0$ (4)
For the case where the generated thinking length exceeds the target ($n_{y}>n_{gold}$), which corresponds to the original LPCO Max Length Constraint Mode as defined in [2], the reward function becomes:
@@
This clipping behavior has implications for training when $\alpha$ is 0.1: if the model’s average generated thinking length is significantly shorter (e.g., more than 15 words below $n_{\text{gold}}$) than high target lengths like 125 or 150 words, it may consistently receive a zero reward component from this length constraint. This could hinder convergence towards such longer budgets, as the model does not receive a continuous gradient “nudge” from the length reward to increase its thinking length into the desired range. While adjusting $\alpha$ to a much lower value could widen the range before clipping, an overly small $\alpha$ might diminish the length reward’s impact, potentially allowing it to be overshadowed by other rewards in the GRPO objective. In experiments, decreasing $\alpha$ was found to be non-effective. It is crucial that the reward function has a slope (before clipping) to ensure effective gradient propagation in GRPO. The referenced Max Length Constraint Mode [2] is designed to apply a soft constraint, gradually penalizing outputs exceeding the target length rather than imposing a hard cutoff, and it incentivizes token efficiency without sacrificing correctness. The $\delta=0.5$ term helps ensure that correct answers with minor budget violations are still preferred over incorrect ones.
@@
Appendix D Example Model Outputs
@@ 1. Sound-generating animated beings with descriptive adjectives:
Appendix E Prompts
E.1 Filtering Prompt
Prompt for Caption Filtering You are a friendly chatbot whose task it is to filter out bad data. You will get a closed caption corresponding to a video clip. Your task is to state whether the caption is a correct subtitle for deaf or hard-of-hearing people. Correct captions in this task are those that correspond to words that could represent an actual sound being made. This could either include a verb that states an impact or sound types or properties like "sound", "noise" or "music". Incorrect closed captions include sentences that someone is saying in the video clip, or sentences that are not related to the video clip at all. All captions are in English. All captions are within curly brackets or square brackets []. Examples of correct captions include: - "(laughs)" or "(laughter)" - "[XBOX SOUND]" - "[chicken bocking imitation]" - "(cereal grains smacking onto wood)" - "(collision)" Examples of incorrect captions include: - "[ transport ]" - "(Wishes are left to wither by time.)" - "(look, I like my nightmareless sleep; I’$l$ play some scary games when I feel too peaceful)" - "[A calm navy color] [TinyTAN character detail]" - "[Haotian Sword Tower]" Is the following caption correct? Please only answer "yes" or "no"
E.2 Caption Generation Prompt
@@
E.3 Prepended Prompt on Model Generation